Supppose we have a binomial experiment in which success is defined to be a particular quality or attribute that interest us.

  • Suppose n =33 and p=0.20. Can we approximate p^ by a normal distribution?why? Np=
  • Nq=
1
Expert's answer
2021-02-24T06:52:22-0500

"\text{for a binomial distribution}"

"\text{mean μ of a binomial }"

"\mu=np"

"\text{standard deviation of a binomial}"

"\sigma=\sqrt{np(1-p}"

"\text{For a normal distribution, μ should be 3 standard deviations away from 0 and n}"

"\mu- 3\sqrt{np(1-p)} > 0 \text{ or}"

"np>9(1-p);(1)"

"\mu+3\sqrt{np(1-p)}<n \text{or}"

"n(1-p)>9p;nq>9p(2)"


"np = 33*0.2=6.6"

"np <9(1-0.2)=7.2"

"nq =33*(1-0.2)=26.4"

"nq>9p=1.8"

"\text{сondition(1) not met}"

"\text{it is not possible to approximate } \hat{p}"

"\text{by a normal distribution}"

Answer:it is not possible to approximate a binomial experiment to a normal distribution

np = 6.6 nq =26.4